commissioning

Document Sample
commissioning Powered By Docstoc
					Locking activities summary
              M.Barsuglia
     locking meeting Jan 15th 2004




                                     1
               N/W cavity experience



• what is done/what is missing?
• what we have learned from n/w cavities?
• problems encountered?
• things not understood?
• hardware/software improvements?
• how we can improve the work/communication in the
group?

                                                     2
 N/W cavity experience


• photodiodes
• lock acquisition
• linear lock
• OMC
• frequency stabilization
• calibration
• simulation


                            3
                                 Photodiodes


• demod phase adjustment: done with the cavity unlocked
                • use the calibration lines?
                • fine phase adjustment needed? Before/during the
                run?
• gain/noise:
                • gain of B1 limited by frequency noise (“horns”?)
                 HF spectrum limited by photodiode noise
                • possible reduction (difficult before the run) of 100Hz-
                1kHz frequency noise before the run



                                                                      4
Signals for lock acquisition

 linearization
                         Linearized error
                              signal


           m

                        No Linearized error
                               signal




                                   5
                          Lock acquistion

 Constraints on the velocity according to the theory:



                   1          λαB
             Δt  B  vMAX  2  33 μm s
Bandwidth
            
            
Force max   v        FMAX
               MAX            10 μm s
            
                      2m



            Gain due to the linearization:    ~ 10

                                                        6
            Lock acquisition statistics
                                      with the linearized error signal

 24 locking events collected locking and delocking the cavity
 23 lock acquisition at the first attempt, only 1 failed locking
attempt




                                                                      7
                 Lock acquisition statistics
 Relative velocity between the mirrors computed for each locking attempt




                                                     2.5 mm/s: mean value
                                                        of the velocity



                                               8 mm/s: maximum velocity for
                      Failed locking
                        attempt               the lock acquisition success

                       v ~ 12.5 μm s


                                               12.5 mm/s: velocity of the
                                              failed event
             8
                                       μm s                                 8
              Lock acquisition-conclusions

• Lock acquistion works without local controls
• speed limit 10 mm/s:
           • limited by the maximum force (increase of ~ 10 of the
           linear zone)
• lock at the first attempt/little energy transfer (speed ~ 2.5 mm/s)
• lock acquisition dynamics understood (theory/simulation)


   • we can work without local controls also with the
   recombined/full ITF? More energy transfer to 0.6 Hz res?
   • we need local control during lock acquistion for the full ITF?


    improvement in the local control filter                          9
                             Linear locking


• Switch between 2 (or more) algorithms ready in the global control,
not tested
               • linearized signal  pure demod signal
               • gain “boost” (add integrator at low frequency)
• transition not needed at this level of sensitivity
• locking accuracy limited by HF noise


        re-write and test the specs for the locking accuracy
        which is the right signal for power fluctuation meas?
         test the transition for GC debug and future
        applications
                                                                  10
                          Locking accuracy
 Dark fringe displacement distribution (Pr_B1p_ACp)
                                                       30 minutes of data

         m




                      RMS
                   3.5  10 10 m




                                                                     11
                           Locking accuracy
Pr_B1p_ACp & Laser freq noise COHE

                                      Coherence up to 10 Hz



                                       Accuracy without the laser
                   COHERENCE            freq noise contribution




    Accuracy ~       10 12 m

                                                               12
                          OMC locking



• Locking, characterization OK
• Transition B1p B1 understood




    Simulation of the OMC with Siesta?




                                         13
      Frequency stabilization
                  2nd step: DC control of the
                  laser frequency




1st step: stabilization “a la TAMA” (and LIGO)
                                                 14
                  Frequency stabilization



• meeting yesterday  minutes by Francois
• 1st step: trials in december: oscillation 170 Hz not still
understood
• Maybe a problem due to the optical transfer function of the MC
(see talk of R.Gouaty in the last C1 meeting)
• Start of a Siesta simulation




                                                               15
170 Hz oscillation




                     16
            Frequency stabilization
                         2nd step: DC control of the
                         laser frequency




     1st step: stabilization “a la TAMA” (and LIGO)


• DC control of the laser freq is really needed ?      17
                   DC control is really needed?
   • 10 mm arm displacement = 1 MHz
   • OMC locking specs = /10000 = 1.5 MHz
         arm displacement (100 s) ~ 10 mm = 1 MHz < specs




Arm displacement
during C1

              ~ 4 mm



                                     ~ 5 min

                                                            18
                   Frequency stabilization

• specification of frequency noise for the ITF full lock acquistion?


Requirement given by the linewidth of the cavity:
       • linewidth of the cavity ~ 10 Hz
       • 10 Hz/10  1 Hz
       • actual RMS noise ~ 30 Hz (C1 data)


Requirement on the speed:
       • equivalent speed given by the frequency fluct <



                                                                   19
                Frequency stabilization


• How to reduce further frequency noise?
             • acoustic mitigation?
             • re-shaping of the curved mirror locking filter




                                                                20
                            Calibration




• commissioning of the laser calibrator
• progress in the calibration procedures




                                           21
                   N/W cavity simulation

• 3-km cavity simulation ready
• used to understand lock acquisition dynamics
• could be used by alignment people
• laser frequency stabilization
• main problem (also for the future):
              • uncertainty of optical parameters
              • validation/speed of the modal simulation


       • Need to talk with suspension and optics people
       • Need to be used more
                                                           22
             Recombined ITF preparation



• Lock acquisition in 2 steps
              • north arm (with B5)
              • mich and west (B2 and B1)
• First SIESTA simulation done
• Now test with the global control algo (almost completed)




                                                             23
Recombined ITF preparation




                         24
                   Recombined ITF signals

Phd 1




Phd 2




        N cav locked   W cav not locked     25
                Recombined ITF signals



• data contaminated by freq noise, and bad alignmend of cavities
and Michelson
• need to realign the ITF and check signals
• Next week, if possible


 validation of B2 signals




                                                             26
                Lock acquisition with trans
                         signals
  • 3 independent steps: n/w/mich
  •First test of north cavity locking with trans signals in nov
  • Bad signals, not understood



2 sec lock




                                                                  27
   We explore also this way ?
                             BS control




• coils for BS have a 45 degrees




• coupling with picth and yaw? Diagonalization needed?
• data from local controls




                                                         28
        Transfer functions measurement




• with the recombined we need to drive 2 mirror at the same time
(0 or p )




• We can use the calibration software ?




                                                            29
                           Review talks



• M.Barsuglia – experience from cavities locking
• L.Barsotti – recombined ITF plans
• F.Bondu – status of frequency stabilization




• we want add something?



                                                   30

				
DOCUMENT INFO